What is the volume occupied by 1 mole of H2O (l) at STP ?

  •  

Grade:11

1 Answers

Arun
4 years ago
Given density =1 gm/ml =1gm /0.001 L
as we know
1 mole of water has 6.022×10^23 molecules
i.e., 18 gm of water has 6.022×10^23 molecules
so 1 gm of water has 6.022×10^23/18 molecules = 0.3345×10^23 molecules
so 0.3345×10^23 molecules in 0.001 L
so 1 molecule will occupy 0.001/0.3345×10^23 L =0.002989×10^-23 L